A current with sinusoidal time-dependence, whose peak value is 6.55 A and frequency is 671 cycles per second, passes through a long straight wire. A rectangular 18.2 cm
Answer #1

first we have to calculate the fem induced on the rectangualr loop:

\epsilon = \frac{d\phi}{dt}

where:

\phi = \int_r^{r + a} Bbdr = \int_r^{r + a} \frac{\mu_0I_0\sin(\omega t)}{2\pi r}bdr =

we havbe the flux as fucntion of time:

\phi = \frac{\mu_0I_0 b\sin(\omega t) }{2\pi}\ln\frac{r +a}{r}

the induced emf is:

\epsilon= \frac{d\phi}{dt}= \frac{\mu_0I_0\omega b\cos(\omega t) }{2\pi}\ln\frac{r +a}{r}

where I0 = 6.55 A w = 2pi f = 671

r = 0.202 m a = 0.182 m and b = 0.0819 m

the power dissipated is:

P = \epsilon^2/R